I have one more question to ask and then I am done on this end for help until I can get a tutor.I appreciate all of your help!
The velocity v (in feet per second) of a rock dropped into a well is related to time (in seconds) by the formula v = 32t. What is the velocity of the rock at the indicated instant? Write your answer in the form (t, v).
(a) when the rock is released ( I amswered on the first part (0,32t) was I correct? I counld not get the answer for the 9 seconds
( , )
(b) after 9 seconds
( , )

Answer by Alan3354(69443)

The velocity v (in feet per second) of a rock dropped into a well is related to time (in seconds) by the formula v = 32t. What is the velocity of the rock at the indicated instant? Write your answer in the form (t, v).
(a) when the rock is released ( I amswered on the first part (0,32t) was I correct?
At release, t = 0
v = 32*0 = 0
--> (0,0)
------------
(b) after 9 seconds
v = 32t = 32*9 = 288 ft/sec
(9,288)